OpenStudy (anonymous):

Solve (x – 5)2 – 100 = 0.

OpenStudy (anonymous):

2x-10=100 2x=110 x=55 Check: (55-5)*2-100=0 50*2-100=0 True

OpenStudy (anonymous):

Based on your other questions, I'm assuming that when you type (x-5)2, rather than meaning \[2(x-5)\] you instead mean \[{(x-5)}^2\] Thus, since your question is \[(x-5)^{2}=100\] The solutions are 15 and -5 (just take the square root of both sides and add 5).

OpenStudy (unklerhaukus):

But that two is a power \[(x – 5)^2 – 100 = 0\]\[(x – 5)^2 = 100 \]\[(x – 5) = 10 \]\[(x – 5) = 10 \]\[x = 15\]

OpenStudy (unklerhaukus):

imperialist know whats what, and didn't leave out the negative solution to the square root

OpenStudy (anonymous):

Thanks for paying attention. I did the simple version :-))) |x-5|=10 x-5=10 and x-5=-10 x=15 and x=-5

OpenStudy (anonymous):

,.ahaha .,happy birthday